# Feature Manipulation for DDPM based Change Detection

### v1

@@ Abstract
Change Detection is a classic task of computer vision that receives a bi-temporal image pair as input and separates the semantically changed and unchanged regions of it. The diffusion model is used in image synthesis and as a feature extractor and has been applied to various downstream tasks. Using this, a feature map is extracted from the pre-trained diffusion model from the large-scale data set, and changes are detected through the additional network. On the one hand, the current diffusion-based change detection approach focuses only on extracting a good feature map using the diffusion model. It obtains and uses differences without further adjustment to the created feature map. Our method focuses on manipulating the feature map extracted from the Diffusion Model to be more semantically useful, and for this, we propose two methods: Feature Attention and FDAF. Our model with Feature Attention achieved a state-of-the-art F1 score (90.18) and IoU (83.86) on the LEVIR-CD dataset.
@@ III Methodology
In examining the baseline model for change detection, we identified two significant limitations impacting its performance. Firstly, the model lacks a mechanism to integrate semantic information within individual temporal branches. Secondly, it fails to capture the temporal correlations between these branches. We introduced two novel feature attention mechanisms into the baseline architecture to mitigate these shortcomings. The enhancements comprise the Siamese Semantic Reasoning (Siam-SR) block [20] and the Cross-temporal Semantic Reasoning (Cot-SR) block [21] are shown on Fig.1a. The Siam-SR block is designed to process and encode semantic information specific to each temporal branch, ensuring that the model recognizes and utilizes the contextual details within individual time points. Concurrently, the Cot-SR block identifies and models the correlations across different temporal branches, facilitating a more cohesive understanding of the changes over time. Fig.2 and Fig.3 depict the architectural integration of these blocks, illustrating how they complement each other to enrich the model’s capacity for nuanced change detection.
@@
In pursuit of this objective, we have elected to implement a strategy delineated by ChangerEx, as depicted in Figure 1b, named the Flow Dual-Alignment Fusion (FDAF). This sophisticated technique engenders the extraction of dual feature sets via the diffusion model [23], which are subsequently subjected to a convolutional network, specifically Flownet, followed by an image warping procedure before their integration into the Change Detection Network. The essence of this approach lies in its utilization of image warping to ameliorate the discrepancies between the paired images, thereby mitigating the impact of noise and enhancing the fidelity of the change detection process.
@@
(a) Attention Based Model.
(b) FDAF Based Model
Figure 1: Structure of the baseline model with attention and FDAF.
@@ IV Experiments
In our comprehensive investigation, we conducted an empirical analysis to ascertain the efficacy of change detection methodologies across two publicly available Remote Sensing datasets, LEVIR-CD and WHU-CD. Our evaluation metrics included the F1 score and the Intersection over Union (IoU) to quantitatively assess the accuracy in identifying altered regions within these datasets as shown in Fig.4. Based on Table-I, within the scope of the LEVIR-CD dataset, the foundational model yielded an F1 score of 90.91 and an IoU of 83.35. Enhancements introduced through the singular application of feature attraction augmented these metrics to 91.52 and 84.36, respectively. Conversely, the isolated deployment of the Flow Dual-Alignment Fusion (FDAF) precipitated a diminishment in performance, registering scores of 82.61 and 70.37. Notably, the concurrent application of both methodologies culminated in a marginal modification of the scores to 91.37 and 84.11, respectively.
Parallel observations were recorded with the WHU-CD dataset, where the baseline configuration delivered an F1 score of 92.65 and an IoU of 86.31. Applying feature attraction in isolation slightly reduced the metrics to 92.40 and 85.87, respectively. In stark contrast, the sole implementation of FDAF significantly eroded performance, with scores plummeting to 76.33 and 61.71. The amalgamation of feature attention and FDAF mechanisms resulted in adjusted scores of 90.27 and 82.27, respectively.
These findings illuminate the intricate dynamics of feature attention and FDAF within the context of change detection in remote sensing [25]. The augmentation of feature attention alone demonstrated a propitious impact on the LEVIR-CD dataset, underscoring its utility in enhancing model performance by intricately capturing and leveraging the temporal correlations between bi-temporal image sequences. Conversely, the integration of the FDAF network, while theoretically posited to ameliorate performance by mitigating noise through alignment and fusion mechanisms, empirically manifested a deleterious effect across both datasets. This dichotomy emphasizes the nuanced and dataset-specific efficacy of these advanced methodological interventions in remote sensing change detection.
Figure 2: Structure of Cot-SR blocks.
Figure 3: Structure of Siam-SR blocks.
LEVIR-CD WHU-CD
Method F1 IoU F1 IoU
88.44 82.17 91.24 84.70
90.18 83.86 91.74 84.93
81.57 69.26 71.82 57.60
Baseline Baseline+Attention Baseline+FDAF Baseline+Attention+FDAF 91.68 82.73 90.27 82.27
TABLE I: The average quantitative results on LEVIR-CD and WHU-CD. All the values are in percentage. Higher values of F1 and IoU indicate good CD performance.
(a) LEVIR, A.
(b) LEVIR, B.
(c) LEVIR, A+B.
(d) WHU, A.
(e) WHU, B.
(f) WHU, A+B.
Figure 4: Qualitative results. A is feature attention, B is FDAF.
To execute our experiments, a singular NVIDIA A6000 Ada Generation GPU was meticulously allocated to ensure uniform computational resources across all trials. The datasets utilized, specifically LEVIR-CD and WHU-CD, underwent preprocessing to standardize the input dimensions to 256x256 pixels, aligning with the established baseline methodology. Within the LEVIR-CD dataset, our training and validation cohorts comprised 7,120 and 2,048 instances, respectively, whereas the WHU-CD dataset was partitioned into 5,952 training and 1,488 validation samples.
To navigate the computational demands of our experiments, we adopted a batch size of eight during the training phase. The optimization of the neural networks was meticulously orchestrated utilizing the Binary Cross-Entropy (BCE) loss function, in conjunction with the AdamW optimizer, a decision underscored by its robustness in handling sparse gradients and regularization. Notably, the diffusion model’s parameters were preserved in a static state throughout the optimization process to isolate the effects of our experimental interventions.
The learning rate was strategically initialized at 1e-5, with a policy of linear decay implemented to gradually reduce this rate to zero as the training epochs progressed. This approach was designed to finely tune the model’s parameters without overshooting the optimal loss landscape. For the LEVIR dataset, the training regimen extended over 120 epochs, whereas for the WHU dataset, a slightly condensed schedule of 80 epochs was deemed sufficient. To ensure the capture of the most efficacious model configuration, we instituted a protocol to record the best validation score after each epoch, thereby enabling the identification and retention of the model iteration exhibiting the highest performance metrics.
@@ V Conclusion
Within the ambit of this discourse, we elucidate a pioneering methodology for change detection, predicated on the nuanced manipulation of feature representations. Predominant paradigms have traditionally leveraged diffusion models as quintessential mechanisms for feature extraction, delineating changed regions within image datasets [26]. Our contemplation on the utility of feature maps has precipitated the inception of two novel augmentative networks: Feature Attention and Flow Dual-Alignment Fusion (FDAF), each tailored to refine the change detection process through strategic feature manipulation.
The Feature Attention mechanism is predicated on the hypothesis that discerning and learning the interrelations between dual feature maps can significantly amplify the efficacy of change detection [27]. This approach has manifested state-of-the-art (SOTA) performance across select datasets, underscoring the viability of correlational learning within feature attention frameworks [28]. Conversely, the FDAF paradigm was envisaged to mitigate environmental noise intrinsic to the conditions under which bi-temporal images are captured, thus enhancing the clarity of change detection. Despite its theoretical promise, empirical analyses have illuminated performance decrement issues with FDAF, ostensibly due to its propensity to filter out ambient noise and obfuscate pivotal feature distinctions between comparative images, thereby attenuating the precision of change detection.
In light of these findings, there lies an imperative for future investigation to refine the FDAF mechanism, ensuring that environmental noise can be effectively excised without compromising the integrity of target feature differences. Our research lays a foundational framework that, with iterative enhancements, could be expansively applied within the domain of change detection, offering robust methodologies for accurately identifying temporal and spatial alterations. The path forward involves ameliorating the identified limitations within the FDAF model and exploring our proposed methods’ adaptability and scalability across diverse change detection applications, heralding a new epoch in the evolution of change detection technologies.

### v2

# Advanced Feature Manipulation for Enhanced Change Detection Leveraging Natural Language Models

@@ Abstract
Change detection is a fundamental task in computer vision that processes a bi-temporal image pair to differentiate between semantically altered and unaltered regions. Large language models (LLMs) have been utilized in various domains for their exceptional feature extraction capabilities and have shown promise in numerous downstream applications. In this study, we harness the power of a pre-trained LLM, extracting feature maps from extensive datasets, and employ an auxiliary network to detect changes. Unlike existing LLM-based change detection methods that solely focus on deriving high-quality feature maps, our approach emphasizes the manipulation of these feature maps to enhance semantic relevance.
@@ III Methodology
Upon evaluating the baseline model for change detection, we discerned two pivotal limitations adversely affecting its efficacy. Firstly, the model does not incorporate a mechanism to integrate semantic information within the individual temporal branches. Secondly, it inadequately captures the temporal correlations between these branches. To address these deficiencies, we introduced two novel feature attention mechanisms into the baseline architecture.
@@
In pursuit of this objective, we have elected to implement a strategy delineated by ChangerEx, as depicted in Figure LABEL:fig:1-b, named the Flow Dual-Alignment Fusion (FDAF). This sophisticated technique engenders the extraction of dual feature sets via the diffusion model [21], which are subsequently subjected to a convolutional network, specifically Flownet, followed by an image warping procedure before their integration into the Change Detection Network. The essence of this approach lies in its utilization of image warping to ameliorate the discrepancies between the paired images, thereby mitigating the impact of noise and enhancing the fidelity of the change detection process.
@@ IV Experiments
TBD
@@ V Conclusion
TBD